given the total fixed cost =$600, variable cost /ynit =$20 and revenue/unit =$40, determine the profit that can be made for a volume of 40
Answer #1

Total revenue=P*Q=40*40=1600

Total cost=TFC+TVC=TFC+AVC*Q=600+20(40)=1400

profit=TR-TC=1600-1400=200

Thus profit at 40unit is $200.